OpenStudy (anonymous):

Verify the identity. cot θ ∙ sec θ = csc θ How do you do that? Am I supposed to just plug this into the calc?

OpenStudy (aum):

cot θ = cos θ / sin θ sec θ = 1 / cos θ Now multiply and simplify

OpenStudy (cwrw238):

convert to sines and cosines

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ \cot θ = \sec θ~\cos θ }\) \(\normalsize\color{blue}{ \sec θ = 1~/~\cos θ }\)

OpenStudy (solomonzelman):

just re-write `cotx ` as `csc x cos x` and re-write `sec x` as `1/ cos x` this is the shortest way.

OpenStudy (solomonzelman):

This is to verify using 1 side only

OpenStudy (solomonzelman):

Need any help, ask !

OpenStudy (anonymous):

When you say rewrite...am I plugging csc x cos x 1/ cos x into the calculator rather than the original equation o.o

OpenStudy (aum):

No calculators involved here. No numbers to plug in. Just some basic definitions of trig functions. Write the left hand side in terms of sine and cosine function mentioned in my first reply. Multiply. Use trig definition again to prove LHS equals RHS.